# Multi-Robot Coordination Induced in Hazardous Environments through an Adversarial Graph-Traversal Game

### v1

@@ Abstract
This paper presents a game theoretic formulation of a graph traversal problem, with applications to robots moving through hazardous environments in the presence of an adversary, as in military and security applications. The blue team of robots moves in an environment modeled by a time-varying graph, attempting to reach some goal with minimum cost, while the red team controls how the graph changes to maximize the cost. The problem is formulated as a stochastic game, so that Nash equilibrium strategies can be computed numerically. Bounds are provided for the game value, with a guarantee that it solves the original problem. Numerical simulations demonstrate the results and the effectiveness of this method, particularly showing the benefit of mixing actions for both players, as well as beneficial coordinated behavior, where blue robots split up and/or synchronize to traverse risky edges.
@@ I Introduction
Consider a scenario where multiple robots must traverse difficult terrain to reach a goal in the presence of an adversary, as shown in Figure 1. The robots wish to reach the goal with minimum risk, and so must plan a path through the environment, while the adversary affects the risk through altering the environment, such as by destroying a bridge, or by positioning its own robots, in military or security scenarios. Alternatively, the adversary may simply represent worst-case natural changes. Such a hazardous environment is often modeled as a graph, which indicates both the different paths robots may take to reach the goal and the risk of those paths through edge weights. For example, probabilistic roadmap planners construct a graph from the environment which is then planned over [1, 2]. To account for a changing environment, one may consider finding the shortest (i. e. minimal risk) path through time-varying graphs, such as in [3] which considers finding the shortest time path when edges have with time-varying delays, and in [4, 5], which both consider time-varying costs and delays on edges. The work in [6] specifically considers planning for a team of robots on a graph, formulating it as an optimization problem where robots can support each other. Another methodology to handle uncertainty in robotics problems is to model them as Markov decision processes (MDPs) [7], assuming a transition function is known. In particular, partially observable MDPs [8] and decentralized MDPs [9, 10] have been applied to robotics. However, while the preceding solutions can account for changes in the environment, they rely on knowledge of how it changes, and so they are not suited for adversarial scenarios and do not indicate how a defender might hinder the intruding robots. Therefore, we turn to game theory for results on adversarial interactions on graphs.
@@
Figure 1: An example scenario motivating our problem. The team of robots intend to reach their destination node 7 with minimal risk, but the terrain is uncertain and time-varying, with the potential for the bridge to be destroyed or trees to be downed on the road. Formulating this as a game can capture the worst case scenario.
@@
Attack graph problems frequently model cyber-security problems as games on graphs [11, 12]. Here, the graph represents the possible avenues of attack, which the defenders can influence. For example, [11] considers the defender adding “honeypots” to the network, modifying the graph nodes to be better defended in a Stackelberg game formulation. Closer to our problem, [12] considers that both players incorporate feedback into heuristic strategies to approximate the Nash equilibria. However, in this formulation, the attacker gains permanent control over the resources represented by nodes, rather than temporarily occupying them as in our problem. Another network security problem over graphs is considered in [13], where an intrusion detection problem is formulated as a stochastic game [14]. A particular type of game more closely related to our problem is barrier coverage, which considers how a defender may place sensors to detect intruding robots, which move through the environment in an attempt to avoid detection [15, 16]. While these two works provide Nash equilibrium results, they are not directly applicable to our problem, because they cannot consider real-time feedback in the defender’s actions.
@@
Therefore, this paper’s contributions are as follows. We formulate the adversarial graph traversal problem as a novel stochastic game, which allows for numerical computation of a mixed Nash equilibrium. We provide theoretical results by bounding the game value with security strategies for both players, and we guarantee that the blue player reaches the goal almost surely under its Nash equilibrium policy. Finally, we demonstrate the preceding theoretical results and show the advantage of mixed strategies and coordinated behavior for multiple robots in numerical examples.
@@ II Problem Formulation
We consider a two-player zero-sum game where the blue player moves its robots through a weighted digraph ${G}(t)$ which is controlled by the red player. Specifically, let ${G}(t)=(\mathcal{V},\mathcal{E},W(t))$, where $\mathcal{V}=\{1,2,\dots N\}$ is the set of nodes for $N\in\mathbb{Z}_{>0}$, $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges, and $W(t)\in\mathbb{R}^{N\times N}$ is the time-varying weighted adjacency matrix, where $W_{i,j}(t)\geq 0$ if $(i,j)\in\mathcal{E}$ and $W_{i,j}(t)=0$ otherwise, for $i,j\in\mathcal{V}$. This indicates that only the weights of the edges are time-varying. The graph belongs to a known set of graphs, ${G}(t)\in\mathcal{G}\triangleq\{{G^{1}},{G^{2}},\dots,{G^{K}}\}$ and, similarly, $W(t)\in\mathcal{W}\triangleq\{W^{1},W^{2},\dots,W^{K}\}$, so that ${G^{k}}=(\mathcal{V},\mathcal{E},W^{k})$ for $k=1,2,\dots,K$, where $K\in\mathbb{Z}_{>0}$ is the number of graphs.
There are $M\in\mathbb{Z}_{>0}$ blue robots and each robot $m$ at time $t$ has position $p_{m}(t)\in\mathcal{V}$. We collect all robot positions together into the vector $p(t)=[p_{1},p_{2},\dots,p_{M}]^{T}\in\mathcal{V}^{M}$. The set of valid actions for the blue player is $A_{\text{blue}}(S)={\overrightarrow{\mathcal{N}}^{\text{out}}_{p_{1}}}\times{\overrightarrow{\mathcal{N}}^{\text{out}}_{p_{2}}}\times\dots\times{\overrightarrow{\mathcal{N}}^{\text{out}}_{p_{M}}}$. Without loss of generality, node $N$ is the goal node, and the game ends when all blue robots reach the goal node ($p(t)=\mathbf{1}_{M}N$). We assume that the goal node $N$ is reachable from every other node in the graph and that it has a self-loop with an edge weight of 00, so that the blue robots can reach the goal from any initial conditions and that any robots which do so can remain there without incurring any additional costs. At each time step, based on its action $a_{\text{blue}}(t)$, the blue player incurs a stage cost of
$C\left(S,a_{\text{blue}}(t)\right)=\sum_{m=1}^{M}W_{p_{m},p_{m}^{+}}(t),$ (1)
@@
The red player’s action is to select the graph at the next time step. Specifically, its available actions at time $t$ are specified by the red player’s action graph $\mathcal{G}^{\text{red}}=(\mathcal{V}^{\text{red}},\mathcal{E}^{\text{red}})$, where $\mathcal{V}^{\text{red}}=\{1,2,\dots,K\}$ are the nodes of this graph where node $k$ corresponds to the position graph ${G^{k}}$, and $\mathcal{E}^{\text{red}}\subset\mathcal{V}^{\text{red}}\times\mathcal{V}^{\text{red}}$ is the set of edges. The action graph is unweighted, directed, and weakly connected, and every node has a self loop. The red player’s action is then $a_{\text{red}}(t)\in{\overrightarrow{\mathcal{N}}^{\text{red}}_{k}}$, where ${\overrightarrow{\mathcal{N}}^{\text{red}}_{k}}$ is the set of out neighbors of $k$ in $\mathcal{G}^{\text{red}}$, and ${G^{k}}={G}(t)$. Therefore, its available actions are determined by the current graph, to model cases where multiple steps may be required to change the environment and changes may be irreversible. See Figure 2 for a representation of $\mathcal{G}^{\text{red}}$. Additionally, the red player has a limited amount of ammo $\alpha(t)$, with $\alpha(0)=\overline{\alpha}\in\mathbb{Z}_{\geq 0}$, which is consumed to change the graph. The red player’s set of valid actions is
@@
Now, we can write the entire game state, containing the positions of the blue robots, the current graph, and the red player’s ammo, as $S(t)=(p(t),\mathcal{G}(t),\alpha(t))\in\mathcal{S}\triangleq\mathcal{V}^{M}\times\mathcal{G}\times\{0,1,\dots,\overline{\alpha}\}$. The game dynamics is
@@
To facilitate describing mixed strategies, we formulate a stochastic game, where the state dynamics is a Markov chain and the transition probabilities are influenced by the actions of both players [14]. The Markov chain is $(\mathcal{S},A_{\text{blue}},A_{\text{red}},P)$ and dynamics is deterministic, such that the transition probabilities $P:\mathcal{S}\times\mathcal{S}\times\mathcal{V}^{M}\times\{1,2,\dots,K\}\rightarrow\{0,1\}$ are in accordance with (II). Mixed policies for the blue player and the red player are defined as mappings $\pi_{\text{blue}}:S\rightarrow\Delta A_{\text{blue}}(s)$, where $\Delta A_{\text{blue}}(s)\subset[0,1]^{|A_{\text{blue}}(s)|}$ is the simplex over $A_{\text{blue}}(s)$, and $\pi_{\text{red}}:S\rightarrow\Delta A_{\text{red}}(s)$, respectively, for $s\in\mathcal{S}$. Now, introducing a discount factor $\gamma\in(0,1]$, we can formally write the expected cost of the game as
@@
The zero-cost self-loop on the goal node ensures no further costs are accrued once all blue robots reach the goal. Although in our original problem $\gamma=1$, choosing $\gamma\in(0,1)$ guarantees that equilibrium policies and a value for the zero-sum stochastic game exist [17, 14]. Specifically, we seek a pair of strategies $\pi_{\text{blue}}^{*}(\cdot),\pi_{\text{red}}^{*}(\cdot)$ and a value $V(s)$ which correspond to a Nash equilibrium, so that
@@
for all $s\in\mathcal{S}$ and all valid mixed strategies $\pi_{\text{blue}},\pi_{\text{red}}$. In terms of our original problem where $\gamma=1$, this Nash equilibrium is an approximate solution, and we will also need to guarantee that the blue robots reach the goal under $\pi_{\text{blue}}^{*}$.
@@ III-A Handling Multiple Blue Robots
Thus far, we have considered that there are $M\geq 1$ blue robots, but here we show how we can consider only a single blue robot, without losing generality. Instead of considering multiple robots moving on the graph ${G}(t)$, we can equivalently consider a single robot moving on a modified joint state graph ${\mathcal{G}^{\prime}}(t)=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{W}^{\prime}(t))$, where each node in ${\mathcal{G}^{\prime}}(t)$ corresponds to a position vector $p(t)$ [18]. Therefore, we have a joint graph set $\overline{\mathcal{G}}^{\prime}=\{{\mathcal{G}^{\prime 1}},\dots,{\mathcal{G}^{\prime K}}\}$, where ${\mathcal{G}^{\prime k}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{W}^{\prime k})$ and each joint weighted adjacency matrix $\mathcal{W}^{\prime k}\in\overline{\mathcal{W}}^{\prime}$ corresponds to weighted adjacency matrix $W^{k}$. To construct this, we assign each position vector $p(t)\in\mathcal{V}^{M}$ to a node $\ell$ in ${\mathcal{G}^{\prime}}$. Then, for each node $\ell\in\mathcal{V}^{\prime}$, we determine its out-neighbors by finding the next position for each action of the blue player, and assign the edge weight in $\mathcal{W}^{\prime k}$ based on the total cost of that action in ${G^{k}}$, for each ${G^{k}}\in\mathcal{G}$. Therefore, we are able to convert any game with $M>1$ on graph ${G}(t)$ to a game with $M=1$ on graph ${\mathcal{G}^{\prime}}(t)$. Finally, because blue robots are identical, we can reduce the size of the joint state graph by considering that robots are indistinguishable. Figure 3 shows the joint state graph for two robots for a simple base graph.
Figure 3: (a) shows a simple example graph, and (b) shows the structure of its joint state graph with two blue robots. The position of each node of the JSG indicates what nodes in the original graph it corresponds to. The markers show a position of two robots in the base graph and the corresponding position in the JSG, and the arrows show an equivalent action in both graphs.
@@ III-B Game Solution and Value Preliminaries
Here we provide some preliminaries to aid in analysis. Define the “highest cost” graph as $\overline{G}=(\mathcal{V},\mathcal{E},\overline{W})$, where $\overline{W}_{ij}\triangleq\max_{k}W^{k}_{ij}$ To simplify notation, let $d_{{G}}(i,j)$, for $i,j\in\mathcal{V}$ denote the weighted distance from node $i$ to node $j$ on graph ${G}$. The following condition on the discount factor $\gamma$ will be needed for certain results:
$\gamma$ $\geq 1-\frac{C_{\text{min}}}{d_{\text{max}}},$
$C_{\text{min}}$ $\triangleq\min_{(i,j)\in\mathcal{E}\setminus(N,N),k\in\{1,\dots,K\}}W^{k}_{i,j}$
$d_{\text{max}}$ $\triangleq\max_{p\in\mathcal{V}}d_{\overline{{G}}}(p,N).\vspace{-25pt}$ (6)
$C_{\text{min}}$ is the minimum possible stage cost that can be achieved without $p(t)=N$, which is strictly positive by assumption. $d_{\text{max}}$ is the maximum distance of any node from the goal on the highest cost graph $\overline{{G}}$. Since the discount factor $\gamma$ is not specified by the problem formulation and the right-hand-side of the inequality in (III-B) is strictly less than $1$, $\gamma\in(0,1)$ can always be chosen so that (III-B) is satisfied.

### v2

# Multi-Robot Coordination Induced in an Adversarial Graph-Traversal Game

@@ Abstract
This paper presents a game theoretic formulation of a graph traversal problem, with applications to robots moving through hazardous environments in the presence of an adversary, as in military and security scenarios. The blue team of robots moves in an environment modeled by a time-varying graph, attempting to reach some goal with minimum cost, while the red team controls how the graph changes to maximize the cost. The problem is formulated as a stochastic game, so that Nash equilibrium strategies can be computed numerically. Bounds are provided for the game value, with a guarantee that it solves the original problem. Numerical simulations demonstrate the results and the effectiveness of this method, particularly showing the benefit of mixing actions for both players, as well as beneficial coordinated behavior, where blue robots split up and/or synchronize to traverse risky edges.
@@ I Introduction
Consider a scenario where multiple robots must traverse difficult terrain to reach a goal in the presence of an adversary, as shown in Figure 1. The robots must plan a path through the environment to minimize risk, while the adversary affects the risk by altering the environment, such as by destroying a bridge or by positioning its own robots, in military or security scenarios.
@@
Such a hazardous environment is often modeled as a graph, which indicates both the different paths robots may take to reach the goal and the risk of those paths through edge weights. For example, probabilistic roadmap planners construct a graph from the environment which is then planned over [1, 2]. For a changing environment, one may consider finding the shortest (i. e. minimal risk) path through time-varying graphs, such as in [3, 4, 5] which find shortest time paths when edges have time-varying delays. The work in [6] specifically considers planning for a team of robots on a graph, formulating it as an optimization problem where robots can support each other.
Another methodology to handle uncertainty in robotics problems is to model them as Markov decision processes (MDPs) [7], assuming a transition function is known. In particular, partially observable MDPs [8] and decentralized MDPs [9, 10] have been applied to robotics. Some works consider risk as simply a cost to minimize, such as [11], which considers path planning of autonomous underwater vehicles, and [12], which considers path planning of unmanned aerial vehicles. Other works consider that the risk is constrained while another cost is minimized, like [13], which considers robotic path planning in a stochastic environment. However, while the preceding solutions can account for changes in the environment, including unpredictable ones, they take a one-sided optimization approach, and so they are not suited for adversarial scenarios. The work [14] includes an estimator of the intentions of other, potentially adversarial vehicles but still approaches it as a one-sided path planning problem. Therefore, we turn to game theory for results on adversarial interactions.
@@
Figure 1: An example scenario motivating our problem. The team of robots intend to reach their destination node 7 with minimal risk, while an adversary controls the condition of the terrain, with the capability to destroy the bridge or down trees on the road.
@@
Game theoretic works for robotic path planning include [15, 16]. In [16], multiple robots with different objectives interact in a continuous environment and attempt to minimize risk, modeled with higher-order moments of the cost. In [15], the robots are explicitly cooperative but have differing information on a partially observable Markov decision process. Closer to our attack/defense scenario, attack graph problems frequently model cyber-security as games on graphs, where the graph represents the possible avenues of attack, which the defenders can influence [17, 18]. For example, [17] considers the defender modifying the graph nodes to be better defended in a Stackelberg game formulation. Closer to our problem, [18] considers feedback heuristic strategies for both players to approximate the Nash equilibrium (NE). However, in this formulation, the attacker gains permanent control over nodes, rather than temporarily occupying them as in our problem.
Another network security problem over graphs is considered in [19], where an intrusion detection problem is formulated as a stochastic game [20]. A particular type of game more closely related to our problem is barrier coverage, which considers how a defender may place sensors to detect intruding robots, which attempt to avoid detection [21, 22]. However, these are not directly applicable to our problem, because they cannot consider real-time feedback in the defender’s actions.
@@
Therefore, this paper’s contributions are as follows. We formulate the adversarial graph traversal problem as a novel stochastic game, which allows for numerical computation of a mixed NE. We provide theoretical results by bounding the game value with security strategies for both players, and we guarantee that the blue player reaches the goal almost surely under its NE policy. Finally, we demonstrate the theoretical results and show the advantage of mixed strategies and coordinated behavior for the robots in numerical examples.
@@ II Problem Formulation
We consider a two-player zero-sum game where the blue player moves its robots through a weighted digraph ${G}(t)$ which is controlled by the red player. Defining $\left[n\right]\triangleq\{1,2,\dots n\}$ for $n\in\mathbb{Z}_{>0}$, let ${G}(t)=(\mathcal{V},\mathcal{E},W(t))$, where $\mathcal{V}=\left[N\right]$ is the set of nodes for $N\in\mathbb{Z}_{>0}$, $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges, and $W(t)\in\mathbb{R}^{N\times N}$ is the time-varying weighted adjacency matrix, where $W_{ij}(t)\geq 0$ if $(i,j)\in\mathcal{E}$ and $W_{ij}(t)=0$ otherwise, for $i,j\in\mathcal{V}$. This indicates that only the edge weights are time-varying. The graph belongs to a known set of graphs, ${G}(t)\in\mathcal{G}\triangleq\{{G^{1}},{G^{2}},\dots,{G^{K}}\}$ and, similarly, $W(t)\in\mathcal{W}\triangleq\{W^{1},W^{2},\dots,W^{K}\}$, so that ${G^{k}}=(\mathcal{V},\mathcal{E},W^{k})$ for $k\in\left[K\right]$, where $K\in\mathbb{Z}_{>0}$ is the number of graphs.
There are $M\in\mathbb{Z}_{>0}$ blue robots and each robot $m$ at time $t$ has position $p_{m}(t)\in\mathcal{V}$. We collect all robot positions into the vector $p(t)=[p_{1},p_{2},\dots,p_{M}]^{T}\in\mathcal{V}^{M}$. The set of valid actions for the blue player is $A_{\text{blue}}(S)={\overrightarrow{\mathcal{N}}^{\text{out}}_{p_{1}}}\times{\overrightarrow{\mathcal{N}}^{\text{out}}_{p_{2}}}\times\dots\times{\overrightarrow{\mathcal{N}}^{\text{out}}_{p_{M}}}$, where $S$ is the game state defined later and ${\overrightarrow{\mathcal{N}}^{\text{out}}_{p}}$ is the set of out neighbors of node $p$. Without loss of generality, node $N$ is the goal node, and the game ends when all blue robots reach the goal node ($p(t)=\mathbf{1}_{M}N$). We assume that the goal node $N$ is reachable from every other node in the graph and that it has a self-loop with an edge weight of 00, so that the blue robots can reach the goal from any initial conditions and that any robots which do so can remain there without incurring any additional costs. At each time step, based on its action $a_{\text{blue}}(t)=p(t+1)$, the blue player incurs a stage cost of
$C\left(S,a_{\text{blue}}(t)\right)=\sum_{m=1}^{M}W_{p_{m}p_{m}^{+}}(t),$ (1)
@@
The red player’s action is to select the graph at the next time step. Specifically, its available actions at time $t$ are specified by the red player’s action graph $\mathcal{G}^{\text{red}}=(\mathcal{V}^{\text{red}},\mathcal{E}^{\text{red}})$, where $\mathcal{V}^{\text{red}}=\left[K\right]$ are the nodes of this graph where node $k$ corresponds to the position graph ${G^{k}}$, and $\mathcal{E}^{\text{red}}\subset\mathcal{V}^{\text{red}}\times\mathcal{V}^{\text{red}}$ is the set of edges. The action graph is unweighted, directed, and weakly connected, and every node has a self loop. The red player’s action is then $a_{\text{red}}(t)\in{\overrightarrow{\mathcal{N}}^{\text{red}}_{k}}$, where ${\overrightarrow{\mathcal{N}}^{\text{red}}_{k}}$ is the set of out neighbors of $k$ in $\mathcal{G}^{\text{red}}$, and ${G^{k}}={G}(t)$. Therefore, its available actions are determined by the current graph, to model cases where multiple steps may be required to change the environment and changes may be irreversible. See Figure 2 for a representation of $\mathcal{G}^{\text{red}}$. Additionally, the red player has a limited amount of ammo $\alpha(t)$, with $\alpha(0)=\overline{\alpha}\in\mathbb{Z}_{\geq 0}$, which is consumed to change the graph. The red player’s set of valid actions is
@@
Now, we can write the entire game state, containing the positions of the blue robots, the current graph, and the red player’s ammo, as $S(t)=(p(t),\mathcal{G}(t),\alpha(t))\in\mathcal{S}\triangleq\mathcal{V}^{M}\times\mathcal{G}\times\left[\overline{\alpha}\right]$. The game dynamics is
@@
To facilitate describing mixed strategies, we formulate a stochastic game, where the state dynamics is a Markov chain and the transition probabilities are influenced by the actions of both players [20]. The Markov chain is $(\mathcal{S},A_{\text{blue}},A_{\text{red}},P)$ and dynamics is deterministic, such that the transition probabilities $P:\mathcal{S}\times\mathcal{S}\times\mathcal{V}^{M}\times\left[K\right]\rightarrow\{0,1\}$ are in accordance with (II). Mixed policies for the blue player and the red player are defined as mappings $\pi_{\text{blue}}:\mathcal{S}\rightarrow\Delta A_{\text{blue}}(S)$, where $\Delta A_{\text{blue}}(S)\subset[0,1]^{|A_{\text{blue}}(S)|}$ is the simplex over $A_{\text{blue}}(S)$, and $\pi_{\text{red}}:\mathcal{S}\rightarrow\Delta A_{\text{red}}(S)$, respectively. Now, introducing a discount factor $\gamma\in(0,1]$, we can formally write the expected cost of the game as
@@
The zero-cost self-loop on the goal node ensures no further costs are accrued once all blue robots reach the goal. Although in our original problem $\gamma=1$, choosing $\gamma\in(0,1)$ guarantees that equilibrium policies and a value for the zero-sum stochastic game exist [23, 20]. Specifically, we seek a pair of strategies $\pi_{\text{blue}}^{*}(\cdot),\pi_{\text{red}}^{*}(\cdot)$ and a value $V(s)$ which correspond to a NE, so that
@@
for all $s\in\mathcal{S}$ and all valid mixed strategies $\pi_{\text{blue}},\pi_{\text{red}}$. In terms of our original problem where $\gamma=1$, this NE is an approximate solution, and we will also need to guarantee that the blue robots reach the goal under $\pi_{\text{blue}}^{*}$.
@@ III Theoretical Analysis
We provide some theoretical guarantees before moving on to the main numerical method. Proofs of theorems are omitted for space and the interested reader is referred to [24].
@@ III-A Handling Multiple Blue Robots
Thus far, we have considered that there are $M\geq 1$ blue robots, but we can consider only a single blue robot, without losing generality. Instead of considering multiple robots moving on the graph ${G}(t)$, we can equivalently consider a single robot moving on a modified joint state graph ${\mathcal{G}^{\prime}}(t)=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{W}^{\prime}(t))$, where each node in ${\mathcal{G}^{\prime}}(t)$ corresponds to a position vector $p(t)$ [25]. Therefore, we have a joint graph set $\overline{\mathcal{G}}^{\prime}=\{{\mathcal{G}^{\prime 1}},\dots,{\mathcal{G}^{\prime K}}\}$, where ${\mathcal{G}^{\prime k}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{W}^{\prime k})$ and each joint weighted adjacency matrix $\mathcal{W}^{\prime k}\in\overline{\mathcal{W}}^{\prime}$ corresponds to weighted adjacency matrix $W^{k}$. To construct this, we assign each position vector $p(t)\in\mathcal{V}^{M}$ to a node $\ell$ in ${\mathcal{G}^{\prime}}$. Then, for each node $\ell\in\mathcal{V}^{\prime}$, we determine its out-neighbors by finding the next position for each action of the blue player, and assign the edge weight in $\mathcal{W}^{\prime k}$ to be the cost of that action in ${G^{k}}$, for each ${G^{k}}\in\mathcal{G}$. Therefore, we are able to convert any game with $M>1$ on graph ${G}(t)$ to a game with $M=1$ on graph ${\mathcal{G}^{\prime}}(t)$. Finally, we can reduce the size of the joint state graph by considering indistinguishable robots. Figure 3 shows an example joint state graph for two robots.
Figure 3: (a) shows a simple example graph, and (b) shows its joint state graph with two robots. The position of each node of the JSG indicates what nodes in the original graph it corresponds to. The markers show equivalent positions and the arrows equivalent actions in both graphs.
@@ III-B Game Solution and Value Preliminaries
Here we provide some preliminaries to aid in analysis. Define the “highest cost” graph as $\overline{G}=(\mathcal{V},\mathcal{E},\overline{W})$, where $\overline{W}_{ij}\triangleq\max_{k}W^{k}_{ij}$ To simplify notation, let $d_{{G}}(i,j)$, for $i,j\in\mathcal{V}$ denote the weighted distance from node $i$ to node $j$ on graph ${G}$. Defining $\mathcal{E}^{-}\triangleq\mathcal{E}\setminus(N,N)$, the following condition on the discount factor $\gamma$ will be useful later:
$\gamma$ $\geq 1-\frac{C_{\text{min}}}{d_{\text{max}}},$ (6)
$C_{\text{min}}$ $\triangleq\min_{(i,j)\in\mathcal{E}^{-},k\in\left[K\right]}W^{k}_{i,j},\quad d_{\text{max}}\triangleq\max_{p\in\mathcal{V}}d_{\overline{{G}}}(p,N).\vspace{-25pt}$
$C_{\text{min}}$ is the minimum possible stage cost that can be achieved without $p(t)=N$, which is strictly positive by assumption. $d_{\text{max}}$ is the maximum distance of any node from the goal on the highest cost graph $\overline{{G}}$. Since the right-hand-side of the inequality in (6) is strictly less than $1$, $\gamma\in(0,1)$ can always be chosen so that (6) is satisfied.